OpenStudy (anonymous):

find dy/dx 4(x-1)^3-2x^2y^3=5x^2

zepdrix (zepdrix):

\[\Large 4(x-1)^3-2x^2y^3\quad=\quad 5x^2\]So we need to find dy/dx implicitly? :) Did I copy the problem down correctly?

zepdrix (zepdrix):

\[\Large \color{royalblue}{\left[4(x-1)^3\right]'}-\color{royalblue}{\left[2x^2\right]'}y^3-2x^2\color{royalblue}{\left[y^3\right]'}\quad=\quad \color{royalblue}{\left[5x^2\right]'}\] So these are the derivatives we'll need to take, in blue. The product rule was applied to the 2x^2y^3 term. Understand that part? :o